OpenStudy (cutiekawdfi):

dentify the conic section, lines of symmetry, and the domain and range. 1. x^2+y2=4 2. x=2y^2+5

OpenStudy (anonymous):

what subject is it trig ?

OpenStudy (anonymous):

idk if this is right but i got (0,0)

OpenStudy (caozeyuan):

for 1 you mean y^2, not y2, right?

OpenStudy (cutiekawdfi):

yes

OpenStudy (caozeyuan):

ok, so it is a circle, x and y axis are symmetry lines

OpenStudy (cutiekawdfi):

okay

OpenStudy (caozeyuan):

2 would be parabola, line of symmetry is y axis

OpenStudy (caozeyuan):

domain for 2 is \[\mathbb{R} \], range is\[\left[ 5,\infty\right]\]

OpenStudy (caozeyuan):

no, I screw up! range is R, domain is 5 to inf

OpenStudy (cutiekawdfi):

:) what about the second one

OpenStudy (caozeyuan):

the second one is done. I left the domain and range of first one for you to figure out, that should be simple enough

OpenStudy (cutiekawdfi):

okay thank you